OpenStudy (anonymous):

A chemist wishes to make 3 liters of a 7% acid solution by mixing a 9% acid solution and a 4% acid solution. How many liters of each solution should the chemist use? Write your answer as a complete sentence.

OpenStudy (nikato):

can you make an equation with this?

OpenStudy (nikato):

its 0.09x+0.04(3-x)=0.07(3) can u solve for x?

OpenStudy (anonymous):

Oh, okay. Thank you both so much!

OpenStudy (anonymous):

First, Lets identify the given Final solution: 3 liters of 7% solution= 0.7(3) Given: 9% acid solution measured in "x" liters. 4% acid solution measured in "x" liters. Solution 0.09x+0.04x=0.7(3) 0.13x=0.27 x=1.615 so the answer is 0.09x=0.09(1.615)=.145 L 0.04x=0.04(1.615)=0.065 L Answer: The acids needed to make a 3 liters 7% acid solution is 0.145 L and 0.065 L solution.

OpenStudy (anonymous):

Welcome c:

OpenStudy (nikato):

so what dud u get as ur answer?

OpenStudy (anonymous):

lol dud :P

OpenStudy (nikato):

wait, i got two different answers. the 9% is 1.8 and the 4% is 1.2